OpenStudy (anonymous):

simplify 18y^3 over 36xy^2 A 1/2x B y/2x C y^2/2x D 1/2xy

OpenStudy (saifoo.khan):

\[\frac{18y^3}{36y^2}= \frac{1y^3}{2y^2}=\frac{1y^{3-2}}{2}=\frac{1y^1}{2}=\Large \frac y2\]

OpenStudy (saifoo.khan):

sorry, that's y/2x

OpenStudy (anonymous):

Answer B

OpenStudy (saifoo.khan):

\[\frac{18y^3}{36xy^2}= \frac{1y^3}{2xy^2}=\frac{1y^{3-2}}{2x}=\frac{1y^1}{2x}=\Large \frac {y}{2x}\]